OpenStudy (anonymous):

Using the given zero, find one other zero of f(x). Explain the process you used to find your solution. 1 - 6i is a zero of f(x) = x^4 - 2x^3 + 38x^2 - 2x + 37.

OpenStudy (anonymous):

Because 1-6i is a zero, 1+6i also must be a 0

OpenStudy (anonymous):

I know how to do the 1-6i part, and I got the x^2+x+37, but how do I find the other roots?

OpenStudy (anonymous):

quadratic

OpenStudy (anonymous):

I know that... But is there any way to find the (x-a) roots?

OpenStudy (anonymous):

no, 37 is prime

OpenStudy (anonymous):

unless you did the 1-6i part wrong lol

OpenStudy (anonymous):

(x-(1-6i)) and (x+(1-6i)) right?

OpenStudy (anonymous):

I mean when you find all the zeros for one of these, you get (x-a), (x-a), (x-(1-6i)), and (x-(1+6i))..

OpenStudy (anonymous):

Im not sure I follow you

OpenStudy (anonymous):

its not necessarily both (x-a)

OpenStudy (anonymous):

I know that.. that's just the formula. Normally when they give you a problem like this, they want you to do the opposite, and they give you all of the roots. For example, earlier I did a problem, and the roots wer (x-3), (x+13), (x+(5+4i)), and (x-(5-4i)). See what I mean?

OpenStudy (anonymous):

@Isaiah.Feynman

OpenStudy (isaiah.feynman):

Shoot! I forget this..

OpenStudy (anonymous):

:c Think harddd.

OpenStudy (anonymous):

here we go again it is actually very easy

OpenStudy (anonymous):

\[5+4i\] is a zero so lets work backwards \[x=5+4i\\ x-5=4i\\ (x-5)^2=(4i)^2=-16\\ x^2-10x+25=-16\\ x^2-10x+41=0\]

OpenStudy (anonymous):

easiest of all is to memorize that if \(a+bi\) is the zero of a quadratic, then it is \[x^2-2ax+(a^2+b^2)=0\]

OpenStudy (anonymous):

in this case you have \(1-6i\) and \(1+6i\) so your quadratic is \[x^2-2x+(1^2+6^2)=x^2-2x+17\]

OpenStudy (anonymous):

oops that was wrong \[x^2-2x+(1^1+6^2)=x^2-2x+37\]

OpenStudy (anonymous):

Why didn't I get the 2x in the middle, or why did you?

OpenStudy (anonymous):

ok i see that you almost had that not quite, but almost so you have \[(x^2-2x+37)(ax^2+bx+c)=x^4 - 2x^3 + 38x^2 - 2x + 37\] and it should be clear that \(a=1\) and \(b=1\) making it \[(x^2-2x+37)(x^2+bc+1)=x^4 - 2x^3 + 38x^2 - 2x + 37\]

OpenStudy (anonymous):

i got it by memorization, but working backwards also give it to you \[x=1+6i\\ x-1=6i\\ (x-1)^2=(6i)^2=-36\\ x^2-2x+1=-36\\ x^2-2x+37=0\]

OpenStudy (anonymous):

assuming you get that, are you good to this step \[(x^2-2x+37)(x^2+bc+1)=x^4 - 2x^3 + 38x^2 - 2x + 37\] ? because we are still not done

OpenStudy (anonymous):

Oh, I see what I did. Nevermind. Continue.

OpenStudy (anonymous):

On the second group of three, how did you know the last number is a 1?

OpenStudy (anonymous):

first of all a made a stupid typo, let me correct it\[(x^2-2x+37)(x^2+bx+1)=x^4 - 2x^3 + 38x^2 - 2x + 37\]

OpenStudy (anonymous):

take a look and see if it is obvious or not that the last number in the second trinomial has to be 1

OpenStudy (anonymous):

Haha. You're good. And I see the 1 now, because we have to keep that 37.

OpenStudy (anonymous):

ok good just like we know the first term is \(x^2\) and not for example \(3x^2\)

OpenStudy (anonymous):

Exactly. Let's continue.

OpenStudy (anonymous):

now how about \(b\) ?

OpenStudy (anonymous):

we have a couple choices lets do an easy one on the right hand side of the equal sign you have \(-2x\) where is the \(x\) term going to come from on the left hand side?

OpenStudy (anonymous):

b would be a 2, right? So that you would multiply it by the x^2, and get 2x^3.

OpenStudy (anonymous):

when you multiply out on the left, there will be two place to get an \(x\) term: \(37\times bx\) and \(1\times (-2x)\)

OpenStudy (anonymous):

that means \(37bx-2x\) has to be \(-2x\) i.e. \[37b-2=-2\] making \(b=0\)

OpenStudy (anonymous):

So lost me there. Hang on...

OpenStudy (anonymous):

there are a couple places on the left to get an \(x^3\) term, not just one

OpenStudy (anonymous):

when you multiply \((x^2-2x+37)(x^2+bx+1)\) there will be two product that will give you \(x^3\) they come form \(x^2\times bx\) and also \(-2x\times x^2\)

OpenStudy (anonymous):

this means \(bx^3-2x^3=-2x^3\) (because that is what you have on the right) so once again \(b=0\)

OpenStudy (anonymous):

Where did you get a -2?

OpenStudy (anonymous):

on the right hand side or on the left hand side? the \(-2x^3\) is on the right hand side because it is

OpenStudy (anonymous):

lets look at this again \[(x^2-2x+37)(x^2+bx+1)\]

OpenStudy (anonymous):

I wrote it down wrong. Ugh. Okay, I'm caught up now.

OpenStudy (anonymous):

ok so you see that if you multiply out on the left, there will be two products to give an \(x^3\) term one is \(x^2\times bx\), the other is \(-2x\times x^2\) when you go to combine like terms you will have \[bx^3-2x^2\] but this must match on the right with \(-2x^3\) making \(b=0\)

OpenStudy (anonymous):

typo again , i meant \(bx^3-2x^3\) sorry

OpenStudy (anonymous):

Yes, so b=0. Got it.

OpenStudy (anonymous):

right, and we didn't have to pick \(x^3\) term to look at, we could have done it as i did at first matching up the \(x\) terms left and right

OpenStudy (anonymous):

way above i wrote \(37bx-2x\) which has to be \(-2x\) once again making \(b=0\)

OpenStudy (anonymous):

you can pick whichever one you like, it has to work with all of them

OpenStudy (anonymous):

Okay, well now we got back to the original problem, but how do we find the zeros?

OpenStudy (anonymous):

now we know \(b=0\) it is easy

OpenStudy (anonymous):

\[(x^2-2x+37)(x^2+1)=0\]

OpenStudy (anonymous):

we already have the zero of \(x^2-2x+37\) as \(1\pm 6i\) now set \[x^2+1=0\] and solve in your head

OpenStudy (anonymous):

Oh, I knew that. Blonde moment. Okay, so they are going to be imaginary numbers.

OpenStudy (anonymous):

yes, two rather famous ones

OpenStudy (anonymous):

So they would be 1+6i, 1-6i, and +i, and -i?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Awesome! Thanks so much! God bless you! :D

OpenStudy (anonymous):

lets check it www.wolframalpha.com/input/?i=+x^4+-+2x^3+%2B+38x^2+-+2x+%2B+37

OpenStudy (anonymous):

yw, my pleasure

OpenStudy (anonymous):

I have that site on my desktop from one of the last times you helped me. :P Lol. It gave me the same, (x^2+1). We're good. :)

OpenStudy (anonymous):

I'll be back tomorrow. I have lots of quizzes to take. Ha. :p